### v1

# Decentralization of Generative AI via Mixture of Experts for Wireless Networks: A Comprehensive Survey

Abstract
Mixture of Experts (MoE) has emerged as a promising paradigm for scaling model capacity while preserving computational efficiency, particularly in large-scale machine learning architectures such as large language models (LLMs). Recent advances in MoE have facilitated its adoption in wireless networks to address the increasing complexity and heterogeneity of modern communication systems. This paper presents a comprehensive survey of the MoE framework in wireless networks, highlighting its potential in optimizing resource efficiency, improving scalability, and enhancing adaptability across diverse network tasks. We first introduce the fundamental concepts of MoE, including various gating mechanisms and the integration with generative AI (GenAI) and reinforcement learning (RL). Subsequently, we discuss the extensive applications of MoE across critical wireless communication scenarios, such as vehicular networks, unmanned aerial vehicles (UAVs), satellite communications, heterogeneous networks, integrated sensing and communication (ISAC), and mobile edge networks. Furthermore, key applications in channel prediction, physical layer signal processing, radio resource management, network optimization, and security are thoroughly examined. Additionally, we present a detailed overview of open-source datasets that are widely used in MoE-based models to support diverse machine learning tasks. Finally, this survey identifies crucial future research directions for MoE, emphasizing the importance of advanced training techniques, resource-aware gating strategies, and deeper integration with emerging 6G technologies.
Index Terms:
Mixtures of experts, wireless networks, generative AI, reinforcement learning, large language model
I Introduction
I-A Background
In recent years, the integration of artificial intelligence (AI) and machine learning technology in wireless communication systems has undergone substantial development, notably characterized by the deployment of large generative AI (GenAI) models like Generative Pre-trained Transformer (GPT) [1, 2, 3]. As the complexity of evolutionary mobile networks escalates and the demand for enhanced wireless performance intensifies, the utilization of large GenAI models has shown significant potential in improving system efficiency and optimizing overall wireless performance [4, 5, 6, 7]. These GenAI models with billions of parameters possess the ability to process tremendous amounts of wireless data and extract intricate patterns of network behavior, which become instrumental in advancing technologies such as channel state information (CSI) prediction [8], orthogonal frequency division multiplexing (OFDM) symbol decoding [9], and transmit power optimization [10]. The employment of GenAI models has effectively addressed the increasing heterogeneity and dynamism of modern networks, facilitating more efficient resource management, capacity expansion, and enhancement of user experiences [11, 12, 13].
However, despite the substantial potential of GenAI models, their large-scale parameter size presents significant challenges for practical deployment in wireless environments. Prominent examples include LLaMA from Meta, which comprises approximately 65 billion parameters [14], PaLM from Google with 540 billion parameters [15], and GPT-4, which exceeds trillions of parameters [16]. Such immense sizes necessitate considerable computational and storage resources, making it difficult for resource-constrained wireless infrastructures such as access points (APs), edge servers, and mobile devices to directly support models of this magnitude [17, 18, 19, 20]. To alleviate the challenge of resource limitations while guaranteeing computational accuracy, distributed deployment and parallel computing strategies have been widely investigated for large-scale GenAI models [21, 22, 23]. Nonetheless, the distributed design often lacks cooperation among devices, leading to limited scalability and inefficiency [24, 25, 26]. Addressing these constraints necessitates a unified framework that combines the strengths of diverse paradigms, emphasizing localized specialization and cooperative mechanisms, especially in heterogeneous and dynamic wireless environments.
TABLE I: Summary of Related Survey
Ref. Description MoE GenAI Wireless Networks
[27] Reviews on the two-decade evolution of MoE, covering foundational concepts, key advancements, and applications in model selection and expert structure.
[28] Traces the development of MoE, highlighting key advances and its role in deep learning, with a focus on expert specialization and optimization techniques.
[29] Categorizes MoE models into mixture of implicitly localised experts (MILE) and mixture of explicitly localised experts (MELE), discussing their advantages, challenges, and applications in regression and classification tasks.
[30] Explores optimization techniques for MoE inference, focusing on architectural, system, and hardware-level approaches to improve efficiency and scalability.
[31] Investigates the transformative impact of MoE and multimodal learning on GenAI models, focusing on the advanced methodologies of MoE in enhancing GenAI.
[32] Introduces the MoE framework, categorizes MoE models, and explores the applications of different gating mechanisms in LLMs.
[33] Reviews techniques for enhancing the efficiency of LLMs, focusing on model-centric, data-centric, and framework-centric methods to reduce resource demands.
[34] Investigates the applications of GenAI in physical layer communications, such as channel estimation and signal classification, and compares GenAI’s benefits over traditional AI models.
[35] Provides a comprehensive history and survey of AI-generated content (AIGC), focusing on GenAI models such as GANs and VAEs, and discusses their potential applications in wireless communication networks.
[21] Surveys the use of LLMs in telecommunications, including network optimization, traffic management, and security, as well as the integration of GenAI in sixth generation mobile networks.
[36] Explores the integration of MoE and GenAI in the Internet of Vehicles (IoVs), focusing on traffic management, autonomous driving, and collaborative decision-making. IoV
This survey Focuses on the efficient training and deployment of MoE frameworks with advanced GenAI approaches for various wireless network scenarios and wireless technologies.
{forest}
[…]
MoE with GenAI [31, 32, 33],
GenAI for wireless Networks [34, 35, 21] , minimum width=7.0cm, minimum width=3.0cm, xshift=-2.0cm, text width=6.0cm, yshift = 0.77cm ] ] ] [Section II. Overview of MoE, minimum width=4.0cm, base=midpoint, yshift = -0.07cm [Definition and Core Components, minimum width=4.0cm, yshift = -0.147cm [Expert network [37, 38, 39, 40], gating mechanism,
and output layer, text width=6.0cm, xshift=-2.0cm, ] ] [Classification of Gating Mechanism, minimum width=4.0cm, [Dense gate [41, 42, 37, 43, 44, 45, 46], sparse gate [47, 48, 49],
hard gate [50, 51], and other gates [52, 53, 54, 55] , text width=6.0cm, xshift=-2.0cm, yshift = 0.14cm] ] [MoE with Generative AI, minimum width=4.0cm,[GANs[56, 57, 58], VAEs [59, 60, 61, 62],
DMs [63, 64, 65], Transformers [66, 67, 68, 69, 70]. , text width=6.0cm, xshift=-2cm, yshift = -0.005cm, base=midpoint, ] ] [MoE with Reinforcement Learning, minimum width=4.0cm, [Value-based RL [71, 72, 73], policy-based RL [74, 75, 76, 77],
multi-agent RL [78, 79, 80], and other RL [81, 82, 83] , text width=6.0cm, xshift=-2.0cm, yshift = 0.14cm] ] ] [Section III.
MoE in Wireless Networks, minimum width=4.0cm, base=midpoint, [Advantages of MoE
in Wireless networks, minimum width=4.0cm, [ Enhance model capacity, computational efficiency,
feature learning, and versatility , text width=6.0cm, xshift=-2cm, yshift = -0.145cm, base=midpoint]] [MoE in Various Wireless
Network Scenarios, minimum width=4.0cm, [Vehicular Networks [84, 85], UAV [86, 87, 88], satellite
[89, 90] and other 6G scenarios [91, 92, 93, 94], text width=6.0cm, xshift=-2cm, yshift = -0.145cm, base=midpoint] ] ] [Section IV. Applications of
MoE in Wireless Networking, minimum width=4.0cm, base=midpoint, [Physical Layer Communication
and Resource Management, minimum width=4.0cm, yshift = 0.142cm, [Channel estimation and prediction [95, 96, 97],
signal processing and communication [98, 99, 100, 61],
radio resource allocation [101, 102, 103] , text width=6.0cm, xshift=-2cm, yshift = 0.00cm, base=midpoint, ]] [Network Optimization
and Security, minimum width=4.0cm, yshift = -0.15cm, [Traffic network management [104, 105, 106],
edge computing and distributed systems [107, 108, 109],
security and anomaly detection [110, 111, 112] , text width=6.0cm, xshift=-2cm, ] ] ] [Section V. Case Study and Dataset, minimum width=4.0cm, [Case Study of MoE with DRL, minimum width=4.0cm, [ MoE is integrated into a diffusion model-based DRL
to demonstrate performance improvements in DRL , text width=6.0cm, xshift=-2cm, base=midpoint, yshift = -0.01cm] ] [Open-Source Datasets, minimum width=4.0cm, [NLP [113, 114, 115], vision [116, 117, 118], multimodal
[119, 120, 121], wireless networks [122, 123, 113, 124] , text width=6.0cm, xshift=-2cm, yshift = 0.14cm ]] ] [Section VI. Future Directions, minimum width=4.0cm, ] [Section VII. Conclusion, minimum width=4.0cm, ] ]
Figure 1: The structure of this survey.
I-B Motivation
The framework of mixture of experts (MoE) has emerged as an effective paradigm to address diverse computational demands and resource constraints for wireless communication systems [27]. The fundamentals of MoE is to decompose a large model into multiple expert modules, each functioning as an independent sub-network specialized for specific inputs [41]. The expert modules are dynamically managed by a gating mechanism that selectively activates only a subset of experts based on the characteristics of the input data, rather than activating all modules simultaneously. Compared with traditional AI, GenAI models require more computational resources for training due to their complex mission objectives and inference procedures. State-of-the-art GenAI models, such as DeepSeek [69], GLaM [48], and Switch Transformer [50], have effectively utilized MoE frameworks to expand model capacity while minimizing computational costs. The dynamic expert selection and sparse activation mechanism of MoE can effectively mitigate computational complexity and improve energy efficiency for GenAI models, enabling resource-constrained wireless infrastructures to enhance scalability and reduce resource consumption in complex wireless task [42].
Another significant advantage of MoE architecture lies in its ability to perform distributed computing and enable collaborative strategies over local experts [94]. The decentralized implementation of MoE experts enables independent processing, inference, and fine-tuning of local data, enhancing specialization by effectively capturing the unique characteristics of personal input [69]. Through using gating mechanisms, the collaborative strategies inherent in the MoE architecture empower edge nodes and wireless devices to collaborate in dynamic and adaptive manner [47]. Moreover, these collaborative weighting and activation mechanisms present extensibility for multimodal and multi-task processing [125, 126, 127]. By integrating multimodal data in wireless environments, such as signal strength, antenna gain, and channel state, MoE can significantly enhance network perception and optimization capabilities based on a shared expert pool to accommodate the complex wireless modalities [55]. For multi-task processing in heterogeneous wireless networks, different tasks often exhibit interdependencies, where tasks such as power control [128] and beamforming [129] aim to enhance signal-to-noise ratio (SNR), and tasks such as spectrum [130] and channel allocation [131] may be constrained by limited resource. With the dynamic expert selection and task-specific delegation, MoE provides a robust solution to the challenges of task conflicts and relationship modeling, achieving significant improvements in terms of efficiency and adaptability for wireless tasks that require simultaneous optimization of multiple metrics [132].
I-C Comparisons with Related Surveys and Contributions
The framework of MoE has demonstrated substantial potential for optimizing wireless resources, improving energy efficiency, and supporting heterogeneous network management. This paper aims to provide a comprehensive survey on the fundamentals of MoE, along with its applications in GenAI and a broad range of wireless scenarios and technologies. Table I presents an in-depth comparison of related surveys, emphasizing MoE, GenAI, and their implementation in wireless networks. These surveys have primarily focused on summarizing the evolution and optimization of MoE. For instance, the authors in [27] reviewed the historical development of MoE, covering fundamental concepts, statistical properties, and practical applications across various domains. The authors in [28] investigated MoE’s core principles, including expert specialization and routing mechanisms with a focus on the field of deep learning. The authors in [29] classified MoE frameworks into mixture of implicitly localized experts (MILE) and mixture of explicitly localized experts (MELE), discussing their applications in regression and classification tasks. Furthermore, the authors in [30] explored the optimization of MoE across system and hardware levels, investigating techniques such as expert selection, load balancing, and distributed computing strategies. For GenAI, existing studies have contributed to enhancing the efficiency of large models with the framework of MoE. For example, the authors in [31] explored the transformative impact of MoE and multimodal learning on GenAI, discussing the role of MoE in improving training and load balancing for GenAI tasks. The authors in [32] and [33] reviewed training techniques for large language models (LLMs), such as fine-tuning, load balancing strategies, gradient routing optimization, and efficient expert pruning.
Additionally, several studies have investigated the integration of GenAI and wireless communication networks. The authors in [34] surveyed the application of GenAI in physical layer communications, underscoring its role in improving channel estimation, equalization, signal classification, and emerging technologies such as reconfigurable intelligent surface (RIS) and beamforming. Similarly, the authors in [35] offered a review of AI-generated content (AIGC), exploring GenAI models and their potential applications in wireless communication networks, while the authors in [21] discussed the implementation of LLMs into telecommunications, focusing on network optimization, traffic management, and security in sixth generation (6G) mobile networks. Furthermore, the authors in [36] incorporated MoE and GenAI models in the Internet of Vehicles (IoV), addressing challenges in traffic management, autonomous driving, and collaborative decision-making for vehicular networks.
However, existing studies lack sufficient investigation into the integrated potential of MoE and advanced GenAI technologies within the broader context of wireless communication systems. This paper fills this gap by comprehensively exploring the diverse application scenarios of MoE in wireless networks, such as vehicular networks [84, 85], unmanned aerial vehicle (UAV) networks [86, 87, 88], satellite networks [89, 90], and other fundamental 6G scenarios [91, 92, 93, 94]. Through in-depth analysis, we demonstrate the capabilities of MoE in solving complex wireless tasks, such as channel estimation [95, 96, 97], physical layer signal processing [98, 99, 100, 61], radio resource allocation [101, 102, 103], traffic management [104, 105, 106], distributed edge computing [107, 108, 109], and network security [110, 111, 112]. The dynamic expert selection and sparse activation mechanisms of MoE have been developed to improve resource utilization, enhance computational efficiency, and ensure adaptability in heterogeneous wireless environments. The main contributions are summarized as follows.
  • We provide a foundational overview of MoE methodologies, covering classical MoE models as well as their integration with GenAI and reinforcement learning (RL), highlighting MoE’s significant capabilities in addressing complex computational demands and optimizing resource utilization.
  • We illustrate the advantages of applying MoE frameworks in heterogeneous and dynamic wireless systems by extensively examining diverse scenarios such as vehicular networks, UAV networks, satellite communications, and emerging 6G scenarios.
  • We thoroughly investigate innovative applications of MoE in wireless technologies, including channel estimation, physical layer signal processing, radio resource allocation, traffic optimization, distributed computing, and network security, demonstrating MoE’s versatility and effectiveness across a broad range of complex tasks.
  • We present a detailed case study that demonstrates the performance advantages of MoE, along with a comprehensive review of publicly available datasets that are widely utilized in MoE-based models to facilitate further research and experimentation in wireless networking.
Figure 2: A comprehensive timeline illustrating the evolution of MoE models, structured across four dimensions: expert networks, gating mechanisms, MoE integrated with GenAI, and MoE applications. The timeline presents diverse expert networks such as FNNs, SVMs, MLPs, CNNs, and self-attention heads, along with the development of dense, sparse, hard, and other variant gating mechanisms. The framework of MoE is widely integrated with GenAI models, such as GANs, VAEs, DMs, and Transformers. Recent applications span NLP, computer vision, multimodal tasks, and wireless networks, demonstrating the versatility of MoE in modern AI and communication systems.
The rest of this paper is organized as follows. Section II provides an overview of MoE, including its components, gating mechanisms, and the integration with GenAI and RL. Section III presents a comprehensive investigation of MoE for diverse wireless networks. Section IV focuses on the practical applications of MoE in wireless technologies. In Section V, we provide a case study on the integration of MoE with an RL task and summarize open-source datasets that support MoE research. Section VI discusses the future research directions. Finally, the paper is concluded in Section VI. The overall structure of this survey is illustrated in Figure 1.
II Overview of Mixture of Experts
Initially introduced in the early 1990s to overcome the limitations of single-architecture models [41], MoE has undergone substantial advancements in modularity and composition, allowing for integration with diverse expert networks, sophisticated gating mechanisms, and recent advanced GenAI models and RL methods. These developments have established MoE as a foundational technique across various domains, including natural language processing (NLP), computer vision, multimodal learning, and wireless networks. Figure 2 presents a comprehensive timeline that traces the evolution of MoE expert networks, gating strategies, and the combination with GenAI, alongside the growing breadth of applications. Building on this foundation, this section presents the definition and core components of MoE, provides an overview of gating mechanisms, and examines its integration with GenAI models and RL methods.
II-A Definition and Core Components of MoE
MoE is a machine learning framework that incorporates multiple specialized models, referred to as experts, each handling a subset of input data within the training process [29]. As illustrated in Figure 3 (a) and (b), compared with classical deep neural networks (DNNs) that directly process an entire dataset, the MoE framework leverages a gating network to allocate dynamically training samples to the most relevant experts [133]. The core components of MoE consist of the expert network, gating scheme, and output layer.
  • Expert Nework: MoE decomposes a model into smaller, more manageable subnetworks, enabling each expert to specialize in distinct data distributions or feature representations. The expert of MoE is typically a trainable network with variations in architecture, depth, or hyperparameters, ranging from conventional machine learning models to advanced DNNs. Examples of expert networks include support vector machines (SVMs) [37], multilayer perceptrons (MLPs) [134], self-attention heads [38], and hybrid structures that integrate feedforward networks (FNNs) with self-attention heads [39]. In more complex scenarios, entire language models can also serve as the expert network within the MoE framework [40].
  • Gating Mechanism: A gating mechanism is responsible for selecting appropriate experts to deal with input samples. Typically implemented as a lightweight model, a gating network computes a probability distribution over the available experts, determining which experts should be activated for model training and inference, thereby reducing computational overhead. Through joint training with expert networks, the gating network continuously improves expert selection, enhancing task specialization and overall model efficiency. Various gating mechanisms have been proposed to optimize expert selection, including dense, sparse, and hard gating, as well as more advanced schemes such as hierarchical gating and multi-gate mechanisms [32].
  • Output Layer: The output layer aggregates the predictions from activated experts and computes the final output as their combination. The most common approach is a weighted sum, where the weights are determined by the gating network to appropriately scale expert contributions based on their relevance to the input [135]. While this method facilitates smooth integration, more advanced implementations incorporate additional approaches to enhance the output quality. For instance, normalization techniques [136] and residual connections [137] are employed to regulate the scale of expert outputs, preventing any single expert from dominating the aggregation process.
Compared to single architectures designed for uniform data, the expert-structured MoE framework improves specialization and training efficiency for diverse data or tasks, particularly in domains requiring high model capacity and adaptability [138].
II-B Classification of Gating Mechanism
The gating mechanism plays a pivotal role in MoE frameworks, serving as the decision maker that dynamically activates the most appropriate experts to process a specific input. We classify the gating mechanism into four types, including dense gating, which assigns weights to all experts; sparse gating, which selects only a subset of experts for enhancing efficiency; hard gating, deterministically allocating each input to one expert; and variant strategies, designed to enhance flexibility and specialization in expert selection.
II-B1 Dense Gating Mechanism
The dense gating mechanism activates all expert networks by assigning probability weights at each iteration. These weights are typically computed using a softmax function, producing a probability distribution over all available experts. The final output is derived as a weighted sum of all expert outputs, ensuring that each expert contributes to the final prediction. Formally, for an MoE framework consisting of $N$ experts $\{E_{1},\ldots,E_{N}\}$, the output of the dense MoE layer is given by
$\mathcal{M}_{\text{dense }}\left(\mathbf{x};\Theta,\left\{\bm{\theta}_{n}\right\}_{n=1}^{N}\right)=\sum_{n=1}^{N}G(\mathbf{x};\Theta)_{n}E_{n}\left(\mathbf{x};\bm{\theta}_{n}\right),$ (1)
where $\mathbf{x}$ denotes the input data and $E_{n}\left(\mathbf{x};\bm{\theta}_{n}\right)$ is the output of the $n$-th expert parameterized by $\bm{\theta}_{n}$. $G(\mathbf{x};\Theta)_{n}$ represents the gating weight for the $n$-th expert, with $\Theta$ denoted as the parameters of the gating network. If the softmax function is used to compute the gating weight, $G(\mathbf{x};\Theta)_{n}$ is expressed as
$G(\mathbf{x};\Theta)_{n}\triangleq\operatorname{Softmax}\left(G(\mathbf{x};\Theta)\right)_{n}=\frac{\exp\left(G(\mathbf{x};\Theta)_{n}\right)}{\sum_{i=1}^{N}\exp\left(G(\mathbf{x};\Theta)_{i}\right)}.$ (2)
Dense gating has been adopted in multiple early investigations [41, 42, 37], and has more recently been applied to large AI models such as LoRAMoE [43], MoLE [44], and DS-MoE [45]. The dense gating mechanism proves effective in tasks demanding comprehensive expert engagement, especially under high uncertainty or diverse feature distributions [46].
II-B2 Sparse Gating Mechanism
While activating all MoE experts enhances model expressiveness, it also incurs significant computational costs, making it impractical for large-scale applications. Sparse gating mechanism uses the Top-K expert selection approach to limit the number of activated experts per input to $K$, tremendously reducing computational overhead [47]. The gating weight of the $n$-th expert $G(\mathbf{x};\Theta)_{n}$ using the Top-K approach is computed as
$\begin{gathered}G(\mathbf{x};\Theta)_{n}=\operatorname{Softmax}\left(\operatorname{Top-K}\left(G(\mathbf{x};\Theta)\right)\right)_{n},\\ \operatorname{Top-K}\left(G(\mathbf{x};\Theta)\right)_{i}=\begin{cases}G(\mathbf{x};\Theta)_{i},&G(\mathbf{x};\Theta)_{i}\text{ is among the }\\ &\text{top-K values of }G(\mathbf{x};\Theta),{}{}{}\\ -\infty,&\text{otherwise. }\end{cases}\end{gathered}$ (3)
Based on the selective activation of sparse gating mechanism, MoE enables the model to scale efficiently while preserving model capacity, which is particularly important in LLMs where computational efficiency is crucial. Examples of sparse gating implementations include GShard [49] and GLaM [48], which significantly improve scalability by activating only a small subset of experts while maintaining high model performance.
II-B3 Hard Gating Mechanism
Unlike soft and sparse gating, the hard gating mechanism deterministically selects a single expert per input based on the highest gating probability. The output weights of the hard gating mechanism are expressed as
$G(\mathbf{x};\Theta)_{n}=\begin{cases}1,&n=\arg\max_{i}G(\mathbf{x};\Theta)_{i},\\ 0,&\text{otherwise. }\end{cases}$ (4)
Consequently, the final output of MoE layer is provided by the output of the highest-weighted expert as
$\begin{split}\mathcal{M}_{\text{hard }}\left(\mathbf{x};\Theta,\left\{\bm{\theta}_{n}\right\}_{n=1}^{N}\right)=E_{i^{*}}\left(\mathbf{x};\bm{\theta}_{i^{*}}\right),\\ i^{*}=\arg\max_{i}G(\mathbf{x};\Theta)_{i}.\quad\quad\quad\end{split}$ (5)
The hard gating mechanism substantially reduces computational overhead, making it particularly suitable for resource-constrained environments. Despite its simplicity, the hard gating mechanism, as utilized in Switch Transformer [50] and DeepSpeed-MoE [139], has demonstrated competitive performance even when activating only one expert per sample. Nonetheless, the hard gating mechanism may confronted with the challenge of imbalanced expert utilization, where certain experts may be over-utilized while others remain underused. Load-balancing techniques [51] and regularization strategies[140] are essential for ensuring more uniform expert utilization. For example, a router Z-loss is introduced to enable a more uniform output proportion of the gating network, thereby enhancing the stability of model training [141]. Given $B$ input data samples, the Z-loss is formulated as
$L_{Z}=\frac{1}{B}\sum_{b=1}^{B}\left(\log\sum_{n=1}^{N}\exp\left(G\left(\mathbf{x}_{b};\Theta\right)_{n}\right)\right)^{2},$ (6)
where $\mathbf{x}_{b}$ is the $b$-th input data and the loss function helps the $G\left(\mathbf{x}_{b};\Theta\right)$ values converge towards a more balanced distribution.
II-B4 Other Variant Gating Mechanisms
Beyond conventional gating mechanisms, several advanced and specialized gating strategies have been developed to further optimize expert selection. Multi-gate MoE (MMoE) frameworks employ multiple gating networks to handle different tasks, enabling task-specific routing that enhances model specialization [55, 53]. The expert networks of MMoE are shared across all tasks, while each task is assigned an independent gating network that separately determines the optimal weighted combination of experts. For an MMoE framework with $T$ gates, the $t$-th output of the MoE layer is computed as
$\mathcal{M}_{\text{multi-gate }}^{t}\left(\mathbf{x};\Theta_{t},\left\{\bm{\theta}_{n}\right\}_{n=1}^{N}\right)=\sum_{n=1}^{N}G_{t}(\mathbf{x};\Theta_{t})_{n}E_{n}\left(\mathbf{x};\bm{\theta}_{n}\right).$ (7)
This framework enables efficient parameter sharing among tasks while allowing task-specific specialization through dedicated gating mechanisms. Meanwhile, hierarchical gating, as a variant of the gating mechanism, introduces multi-level decision-making, where experts are partitioned into subgroups, and the gating functions are performed in multiple stages to refine expert selection [52, 54]. The output of a two-level hierarchical gating mechanism is expressed as
$\begin{split}&\mathcal{W}_{\text{hier}}\left(\mathbf{x};\Theta^{high},\left\{\Theta_{m}^{low}\right\}_{m=1}^{M},\left\{\bm{\theta}_{m,n}\right\}_{n=1}^{N_{m}}\right)=\\ &~{}\sum_{m=1}^{M}G^{high}(\mathbf{x};\Theta^{high})_{m}\sum_{n=1}^{N_{m}}G_{m}^{low}(\mathbf{x};\Theta_{m}^{low})_{n}E_{m,n}\left(\mathbf{x};\bm{\theta}_{m,n}\right),\end{split}$ (8)
where the high-level gating $G^{high}(\mathbf{x};\Theta^{high})$ first selects expert group $m$ from $M$ available groups for the input $\mathbf{x}$. Within the selected group $m$, the low-level gating $G_{m}^{low}(\mathbf{x};\Theta_{m}^{low})$ further weights the expert output among $N_{m}$ experts, with the $n$-th expert output denoted as $E_{m,n}\left(\mathbf{x};\bm{\theta}_{m,n}\right)$. The hierarchical gating mechanism is well-suited for tasks requiring multi-level decision-making and is particularly effective in wireless network scenarios with strong task correlations.
Lessons Learned: Gating mechanisms are a critical component of the MoE framework, responsible for regulating expert activation and managing the allocation of computational resources. The choice among dense, sparse, hard, and more advanced adaptive gating strategies involves a fundamental trade-off between expressiveness, efficiency, and specialization. Specifically, the dense gating mechanism ensures full expert participation but incurs substantial computational overhead. In contrast, sparse and hard gating mechanims improve efficiency by limiting expert activation, while potentially posing challenges related to stability and capacity utilization. Emerging designs such as multi-gate and hierarchical gating mechanism provide enhanced flexibility by enabling task-aware or multi-tier expert selection, thereby supporting fine-grained specialization. The effectiveness of a gating mechanism depends on balancing computational cost and model capacity, as well as adapting to task complexity and data heterogeneity. As such, the design and deployment of gating mechanisms are pivotal for scaling MoE frameworks to meet the diverse demands of machine learning applications.
Figure 3: Illustration of three learning architectures. (a) Classical DNN model: A straightforward architecture for uniform data distribution but limited by low computational and energy efficiency. (b) MoE framework: Incorporate multiple expert networks coordinated by a gating mechanism for diverse data or tasks. (c) MoE with GenAI: Combine the specialization of MoE with generative capability, illustrated by an example of MoE integrated with multiple encoder experts of variational autoencoder (VAE) [61]. (d) MoE with RL: Employ multiple actor experts (AEs) and critic experts (CEs) to enhance the training performance of the actor-critic (AC) algorithm [73].
II-C MoE with Generative AI
TABLE II: Summary of Generative AI model with Mixture of Experts Integration.
Note: Red circles indicate the limitations of conventional GenAI and green circles highlight the advantages of MoE
Model Ref. Task                   Insight GenAI Limits & MoE Advantages
    GANs [57] Language Generation Improve representation and coherence via expert generators and feature alignment Mode collapse, limited sample diversity and difficulties in multimodal distributions Multi GAN generators improve generation diversity and captures distinct data modes Generator selection enhances semantic alignment and generation efficiency
[57]   Image   Generation Use gating network to dynamically select specialized generators
[58] Text-to-image Generation Apply sparse gating and FNN experts for semantically aligned feature synthesis
    VAEs [60] Anomaly Detection Use convolutional VAEs for manifold-specific modeling Struggle with complex representation in the latent space Enable expert specialization through multiple encoding and decoding modules Enhance model expressiveness by mitigating latent discontinuities
[61] Channel Coding Multi encoders to handle representation discontinuities under bandwidth limits
[62] Multi-modal Recommendation Fuse modality-specific experts based on uncertainty-aware stochastic sampling
  Diffusion   Models [63] Text-to-image Generation Train specialized denoisers for different denoising intervals to align text and image fidelity Fixed parameters across denoising steps hinder adaptability Employ specialized experts for different denoising stages, adapting to distinct noise characteristics at each time step Sparse experts reduce inference cost while maintaining generation quality
[64]  Image  Generation Assign timestep intervals to pruned experts for reducing inference overhead
[65]  Image  Generation Employ timestep-aware MoE for adaptive routing and improved convergence
Transformers [68] Language Generation Construct attention and MLP-based MoE for instructed LLMs with post-training Substantial computational requirements and limited adaptability to diverse modalitie Replace FFNs with MoE layers to reduce computation and maintain model capacity Enable experts to handle inputs across language, vision, and multimodal domain
[69] Language Generation Employ expert segmentation and shared experts to maximize specialization
[70] Multimodal Modeling Combine modality-specific MoE in vision and language tasks
The employment of MoE framework in GenAI models has garnered increasing attention, leveraging MoE experts with generative capabilities to enhance data synthesis, representation learning, and adaptive processing. Recent advancements have demonstrated the effectiveness of MoE in improving the performance of GenAI models, including generative adversarial networks (GANs), variational autoencoders (VAEs), diffusion models (DMs), and Transformer-based architectures [79]. This section explores the integration of MoE and GenAI, highlighting key advancements in optimizing generative performance across diverse applications. Figure 3(c) illustrates an MoE-enhanced GenAI framework based on VAE, and a summary of related approaches is provided in Table II-C.
II-C1 Generative Adversarial Networks
GANs are widely used for high-quality image synthesis, consisting of a generator that synthesizes realistic data samples and a discriminator that differentiates between real and synthetic inputs [142]. Despite their effectiveness in generation capability, conventional GANs often suffer from mode collapse, limited sample diversity, and difficulties in modeling complex multimodal distributions [143]. The integration of MoE into GANs provides a scalable solution by incorporating multiple specialized generators or expert modules within the generator [144]. The gating network in MoE dynamically selects the most relevant expert, improving mode diversity, mitigating mode collapse, and enhancing generalization across heterogeneous data distributions. Recent advancements have explored the integration of MoE framework into GANs to enhance generative modeling across different domains. In the context of language generation, MoE-GAN [56] incorporates multiple expert generators for language feature statistics alignment, which offers fine-grained learning fidelity and improves the diversity and coherence of generated text. Additionally, MEGAN [57] applies MoE to image generation by employing a set of generators and a gating network adaptively selecting suitable generators for each input, thereby capturing distinct data modalities. Extending this concept to text-to-image synthesis, a sparse MoE-based GAN framework is introduced to leverage both textual embeddings and latent variables for specialized feature generation, thus providing more specialized and semantically aligned images [58].
II-C2 Variational Autoencoders
VAEs are probabilistic generative models that encode input data into a latent distribution and reconstruct data from a sampled latent variables [59]. While the encoding-decoding structure of VAEs is effective in data compression and generation, conventional VAEs often struggle with capturing complex representation in the latent space. Exploiting MoE in VAEs enables specialized latent space modeling by dynamically allocating data samples to different encoder or decoder experts, improving reconstruction capability and representation diversity. For instance, in anomaly detection, employing a set of convolutional VAEs as experts, together with a convolutional gating network to select experts for modeling manifold-specific patterns, has been shown to improve detection accuracy [60]. In communication scenarios, an introduction of multiple VAE encoders to mitigate representation discontinuities under bandwidth constraints has led to improved robustness across varying channel conditions [61]. Furthermore, in multi-modal recommendation, a mixture of variational stochastic auto-encoders [62] is leveraged to select and fuse modality-specific latent experts, which effectively enhances generalization for different modalities under incomplete input conditions.
II-C3 Diffusion Models
DMs comprise a forward process, where random noise is gradually added to the original data, and a reverse process that reconstructs structured data through a multi-step denoising process [145]. Conventional DMs use a single set of parameters across all denoising steps, limiting their adaptability to different noise levels [146]. MoE-based diffusion models overcome this constraint by employing specialized experts for different denoising stages based on distinct noise characteristics [147]. Additionally, MoE enhances computational efficiency through sparse expert activation, reducing the need for full-network inference at every denoising step [148]. Recent developments have explored the integration of MoE into diffusion models to improve both generation quality and computational efficiency. For instance, eDiff-I, a text-to-image diffusion model proposed by NVIDIA, has introduced an ensemble of expert denoisers for different generation stages, thereby improving text alignment and image fidelity in text-to-image synthesis [63]. Meanwhile, DiffPruning [64] reduces the inference cost of image generation by assigning clustered timestep intervals to pruned expert networks, while Switch-DiT [65] integrates timestep-specific characteristics within sparse MoE layer to accelerate convergence.
II-C4 Transformers
Transformers have emerged as the prevalent architecture in GenAI models, demonstrating remarkable capabilities in sequential modeling through self-attention mechanisms and encoder-decoder architectures [149]. However, as the model sizes continue to grow, the computational demands of Transformers increase substantially, posing challenges for efficient training and deployment, particularly in LLM applications [150]. A widely adopted solution is to replace the standard feedforward networks (FFNs) in Transformer blocks with MoE layers, where a gating mechanism dynamically selects a sparse subset of smaller FNN for each input token, thereby reducing computation while maintaining model capacity [66]. Moreover, MoE also facilitates expert specialization by enabling different experts to handle distinct input patterns or tasks, enhancing performance across language, vision, and multimodal domains [67]. The effectiveness of MoE-based Transformers has been validated in recent models such as LLaMA-MoE [68] for scalable language modeling, DeepSeek-MoE [69] for cross-layer expert allocation, and MoME [70] for flexible multimodal integration.
Lessons Learned: The integration of MoE into GenAI models has demonstrated significant improvements in computational efficiency, representation learning, and adaptive specialization throughout the generative process. A key insight is that generative models often involve highly structured procedures, such as multi-step denoising [63, 64, 65], sequential decision-making [151, 152], or modality-conditioned synthesis [58, 67]. In these settings, MoE allows GenAI models to activate only the most relevant experts at each stage, enabling fine-grained refinement of the generation process and adaptive utilization of model capacity. The resulting specialization improves fidelity and stability without increasing computational complexity, particularly in models designed to handle heterogeneous data types Furthermore, with the emergence of new generative paradigms such as Generative Flow Networks (GFlowNets) [153], MoE provides a scalable and modular framework that supports expert specialization, enhances generation quality, and addresses the increasing demands of advanced GenAI models.
II-D MoE with Reinforcement Learning
TABLE III: Summary of Reinforcement Learning Methods with Mixture of Experts Integration. Note: Red circles indicate the limitations of conventional RL and green circles highlight the advantages of MoE
[…]
C4 [114] A large-scale corpus derived from web pages and filtered for cleaned content 750GB of clean English text MC-MoE LLM [236]
Question and Answer TeleQnA [115] Knowledge of LLM in lexicon, research overview, publications, standards overview, and specifications 10,000 questions distributed across 5 categories ETR-MoE [237]
{tblr}
[…]
{tblr}
[…]
{tblr}
[…]
[r=1]c Vehicular Network CitySim [124] A high-precision vehicle trajectory dataset over diverse road scenarios, designed for intelligent Internet of Vehicle research. 25,000+ vehicle trajectories covering highway, urban, and intersection scenes MoE-based vehicle trajectory prediction [177]
In the domain of multimodal research, particularly in image-text and video-text generation tasks, datasets such as MS COCO [119], LAION-5B [120], and HowTo100M [121] have been widely adopted to align visual and textual modalities. Specifically, MS COCO provides 330,000 images, each annotated with five human-written captions, which is commonly applied for fine-grained analysis of visual scenes in tandem with language annotations [126]. LAION-5B is an open-source large-scale dataset containing 5.85 billion CLIP-filtered image-text pairs [135] across multiple languages, offering unprecedented scale and diversity for training and evaluating multimodal models in tasks such as retrieval, captioning, and zero-shot classification [242]. Meanwhile, HowTo100M contains 1.22 million narrated instructional videos with 136 million CLIP-caption pairs and automatic speech recognition (ASR) transcripts, enabling multimodal learning for video-text alignment across 23,000+ task types. Multimodal datasets often present challenges in aligning modalities with different sequence lengths, such as text-image or text-video pairs [246]. MoE-based models have demonstrated the ability to perform robust cross-modal matching strategies through specialized experts trained on these multimodal datasets, enabling effective handling of text-image or text-video alignment within a unified framework [243].
V-B4 Dataset for Wireless Networks
Datasets for wireless networks such as RadioML [122], DeepMIMO [123], 5G-NIDD [232], and CitySim [124] serve diverse wireless scenarios ranging from physical-layer signal processing to system-level intrusion detection and vehicular networking. In particular, RadioML offers labeled in-phase/quadrature (IQ) signal recordings, which are utilized by the MoE-based automatic modulation classification (AMC) model to capture distinctive signal features under varying SINR scenarios [244]. Meanwhile, DeepMIMO, built on accurate ray-tracing simulations, provides realistic mmWave and massive MIMO channels tailored to beam selection, channel estimation, and large-antenna array optimizations [245]. For network security, 5G-NIDD presents comprehensive network intrusion detection logs with fully labeled traffic traces derived from operational real-world 5G testbeds. This dataset allows MoE-based intrusion detection systems to differentiate between benign connections, newly emerging threats, and zero-day exploits, reducing false alarms and improving detection agility [111]. Additionally, the CitySim dataset focuses on connected and autonomous vehicles, delivering drone-captured vehicular trajectories that include complex mobility patterns, thereby supporting safety-oriented and MoE-based wireless research for connected and intelligent transportation systems [177].
VI Future Research Directions
MoE has demonstrated promising advances across various wireless scenarios. Nonetheless, several open research directions remain to be explored for fully utilizing its capabilities. This section outlines key future directions, categorized into improving existing solutions and exploring new methodologies to extend MoE’s capabilities in intelligent wireless systems.
VI-A Improving Existing Solutions
VI-A1 Lightweight and Resource-Efficient MoE Architectures
While MoE frameworks offer substantial capacity and adaptability, the intensive computation and memory demands remain challenging for edge devices, such as IoT nodes and UAVs that are power or storage constrained. To overcome these limitations, parameter-reduction strategies, such as expert pruning [247], quantization-aware training [248], and knowledge distillation [249], present promising solutions to selectively retain critical expert network weights, thereby reducing the computational overhead. In addition, collaborative model execution across edge devices and cloud servers offers a compelling direction, where lightweight expert modules are deployed locally for low-latency processing, and more complex components are offloaded to centralized infrastructure.
VI-A2 Dynamic Gating and Load Balancing Scheme
In dynamic wireless environments, sparse gating mechanisms are commonly employed to activate a minimal subset of experts However, sparse gating mechanisms are often susceptible to expert imbalance, where certain experts are disproportionately activated while others remain underutilized. This uneven utilization can lead to suboptimal resource efficiency and performance bottlenecks. To address this challenge, techniques such as load-balancing regularization, entropy-based sparsity constraints, and dynamic gating mechanisms (e.g., multi-objective optimization [250]) can be incorporated to promote more equitable expert allocation. These approaches strike a balance between model capability and resource utilization, enabling more effective MoE operation and sustaining real-time performance in heterogeneous wireless systems.
VI-A3 Communication-Efficient MoE in Distributed Wireless Networks
In distributed wireless networks, the deployment of MoE architectures across multiple edge devices and network nodes introduces significant communication overhead and transmission delays. Efficient coordination among geographically distributed experts is critical to minimize network traffic burden, synchronize communication delays, and reduce energy consumption, especially in bandwidth-constrained wireless edge environments. Investigating efficient task partitioning, asynchronous communication protocols, and data compression and offloading techniques is essential for enabling rapid and lightweight inter-node communication. By optimizing the communication link quality and reducing redundant data flows, these approaches are capable of maintaining robust training and inference of MoE while ensuring real-time performance.
VI-A4 Multimodal MoE for Integrated Wireless Networks
While MoE frameworks have exhibited remarkable performance in multimodal learning tasks involving image-text or video-text data, their application in wireless networks has predominantly remained within unimodal scenarios, typically confined to tasks such as radio frequency signal processing [244]. By incorporating auxiliary data sources such as environmental imagery, inertial sensor measurements, or high-level semantic user intents, expert networks can be designed to extract complementary features across distinct data patterns. In parallel, adaptive gating mechanisms can enable expert selection based on different modality availability, thereby supporting more accurate channel estimation, adaptive beamforming, and interference suppression. Wireless multimodal synthesis holds significant potential to enhance situational awareness and advance the development of communication networks.
VI-B Exploring New Methodologies
VI-B1 Cross-Layer Coordination and Optimization
Current utilization of MoE in wireless protocol stacks conventionally focuses on isolated protocol layers, such as physical‑layer communication or MAC‑layer access control. A promising research direction lies in developing cross‑layer MoE architectures that jointly optimize different layer protocol tasks, including link adaptation, congestion control, traffic scheduling, and routing strategy, within a unified decision‑making framework. Such cross‑layer integration facilitates holistic optimization and ensures that decisions made at individual layers align with global system goals, enhancing overall performance such as spectral efficiency, latency reduction, and reliability.
VI-B2 Hardware Acceleration and Deployment of MoE
Beyond the domain of algorithmic design, advancing MoE for wireless applications requires addressing challenges related to hardware compatibility and system‑level integration. Exploring co‑optimized hardware architectures, such as AI accelerators with embedded MoE gating logic [107] or CPU/GPU design pipelines tailored for sparse computation [69], provides efficient support for expert selection and conditional MoE execution. Scalable deployment of MoE under such conditions demands intelligent runtime orchestration frameworks capable of adapting to heterogeneous hardware capabilities, dynamic network topologies, and evolving wireless infrastructures.
VI-B3 Novel Paradigms in IoT, Digital Twins, and Beyond
As wireless networks evolve toward emerging paradigms such as digital twins, cognitive IoT, and zero‑touch network management [251], MoE frameworks are envisioned to serve as a fundamental approach for dynamic adaptability and low‑latency responsiveness. By tailoring gating mechanisms to contextual factors such as user mobility, service‑level agreements, and energy‑efficiency constraints, next‑generation wireless systems can more effectively coordinate heterogeneous operational objectives with minimal resource overhead, thereby advancing toward context‑aware and data‑driven network intelligence.
VII Conclusion
This survey has comprehensively reviewed the integration of MoE in wireless networks, demonstrating its substantial advantage in enhancing adaptability and improving efficiency across various wireless scenarios. By systematically exploring fundamental MoE methodologies, including various gating mechanisms and their integration with GenAI and RL, the survey has extensively explored applications across key wireless domains such as vehicular networks, UAVs, satellite networks, HetNets, ISAC, and mobile edge computing. Furthermore, critical wireless network tasks, including physical layer communications, radio resource management, network optimization, and security, have been thoroughly discussed. A case study integrating MoE into a diffusion-based DRL framework has revealed its empirical advantages in wireless optimization tasks. Additionally, an overview of open-source datasets has been presented to support ongoing MoE-related experimentation. Future directions should focus on lightweight architectures, balanced expert allocation, and integration with emerging paradigms, facilitating MoE as a pivotal component for next-generation wireless systems.